OpenStudy (anonymous):

solve the following trigonometric equation for the given domain: sin(x)=cos(x) on[0,2π]

OpenStudy (anonymous):

pi/4

OpenStudy (anonymous):

the rest of the solutions would be pi+pi/4, pi-pi/4, 2pi-pi/4

OpenStudy (anonymous):

because remember \[\sin(\pi/4)=\frac{ \sqrt(2) }{ 2 } and \cos(\pi/4)=\frac{ \sqrt(2) }{ 2 } \]

OpenStudy (anonymous):

 then you wanna find all the reference angles.

OpenStudy (anonymous):

Thank you!!!!!!!